# Voyaging into Unbounded Dynamic Scenes from a Single View

### v1

@@ 3.3 Dynamic Scene Update
Point cloud merging. We exploit the depth model [37] to obtain the video depth maps ${\bm{D}}^{(1)}$ of the outpainted video ${\bm{I}}^{(1)}$. Noting that there is a depth inconsistent problem between the estimated depths and the rendered ray depth map ${\bm{\hat{D}}}^{(1)}$ in the mask regions ${\bm{\hat{M}}}^{(1)}$, we follow common practices [58, 57] to align the video depth maps ${\bm{D}}^{(1)}$ with the ray depth maps ${\bm{\hat{D}}}^{(1)}$ by finetuning the depth estimation model. After obtaining the aligned depth maps from the finetuned depth model, we follow §3.1 to initialize foreground point cloud ${\mathcal{P}}^{\mathrm{(1,f)}}$ and background point cloud ${\mathcal{P}}^{\mathrm{(1,b)}}$. Then we merge the new point clouds with the previous one to update the dynamic scenes,
@@
where ${\mathcal{P}}^{(1)}$ is the updated point cloud at the pose $\bm{\Pi}^{(1)}$. In this way, we could generate unbounded dynamic scenes along fly-through camera trajectories by looping §3.2 and §3.3 with new camera poses and scene prompts.
Background completion. In practice, we find that the missing background point clouds occluded by the moving foreground in the scene cause inconsistency when rendering along fly-through camera trajectories (as shown in Figure 4). We hence propose to employ our video outpainting model to generate background videos of the occlusion parts. After that, we extract depth maps of the background videos and reconstruct the background point clouds with the extracted depth maps. Details are in supplementary material. We also update the scene with these point clouds.
@@ 4 Experiments
In this section, we first describe the details of our experiment setups. Then, we discuss the results of the scene generation with fly-through cameras and controllable scene generation. We finally conduct ablation studies of our model.
@@ 4.1 Experiment Setups
We trained our outpainting model with the OpenVid dataset [33], which contains high-resolution videos with detailed video prompts. In order to train with daily dynamic scene videos, we sampled $5\times 10^{3}$ videos that contain outdoor dynamic scenes, including natural scenes (such as waterfalls, rivers and moving clouds) and urban scenes (such as walking people and moving cars). Further data curation details are in supplementary material.
@@
Training details. For training the proposed video outpainting model, we employed the LoRA [18] strategy to finetune the CogvideoX-5B-I2V diffusion model [54] with ControlNet [59]. The LoRA rank was $256$ and the learning rate was $1\times 10^{-4}$ with a cosine schedule. The video resolution is 16$\times$512$\times$512 and hence $N=16$ and $h=w=512$. For calculating the ray information, we introduced the depth model [37] to extract the depth map of each video frame. Then we followed MOTIA [46] and exploited the same mask strategy to generate partial videos.
Camera details. To set up dynamic scenes with fly-through camera trajectories, we followed the common practices [58, 35, 23] to move cameras following a straight line or doing a rotation. For the straight line, we used 0.0005 camera speed to move backward. For camera rotation, the rotation radians were 0.45. We generated camera paths by linearly interpolating the translation and rotation between two camera poses. Along each path, we evenly sampled 16 views to synthesize fly-through videos.
@@ 4.2 Scene Generation with Fly-through Cameras
Figure 6 shows the dynamic scene generation results with long fly-through trajectories. We tested our model on both the background dynamics (water flowing in the waterfall) and the foreground dynamics (person walking on the path). Compared to the initial point cloud, our model generated unbounded dynamic scenes by exploiting the proposed scene outpainting process. As there is no direct baseline for unbounded dynamic scene generation, we compare our method with unbounded static scene generation method WonderJourney [58] in Figure 7, state-of-the-art  4D scene generation model CAT4D [49] in Figure 2 and camera-controllable video diffusion model MotionCtrl [48] in Figure 8. These figures demonstrate that our model can better generate unbounded dynamic scenes with 3D consistent motions by exploiting our ray outpainting model.
@@ 4.3 Scene Generation Controlled by Text
To test the controllable dynamic scene generation ability of our model, we generated dynamic scenes from village and volcano images with scene prompts. As shown in Figure 9, for the village image, we control the outpainting content by introducing two different text prompts: waterfall and huge lake. For the volcano image, we prompted the model to generate glacier and magma. It can be seen that our model successfully controls the unbounded dynamic scene generation with the given scene prompts. We followed the common practices [56, 23, 58, 57] to employ the CLIP [36] scores between the novel view images and the input scene prompts and the dynamic degree (DD) and factual consistency (FC) indices proposed by VideoScore [14] to thoroughly evaluate the generation quality of the video. Table 1 shows that, while WonderJourney only generates static scenes with scene prompts, our model achieves better performance by controlling dynamic scene generation.
@@ 4.4 Ablation Study
Ray outpainting. To fairly evaluate the proposed ray outpainting model, we conducted experiments on five scenes containing common dynamic objects: cars, people, water and clouds. For each scene, we sampled 16 novel view images from the generated dynamic point cloud at 16 different timestamps along the fly-through camera trajectories. As there is no 3D video outpainting model baselines, we chose 2D outpainting model MOTIA [46] for comparisons. Figure 5 shows that by introducing the ray information for outpainting, our model successfully outpaints 3D consistent motions at the outpainting boundaries. In this way, our ray outpainting model achieves better performance in Table 2.
@@ 5 Conclusion
In this paper, we propose DynamicVoyager to voyage into unbounded dynamic scenes from a single view. We reformulate dynamic scene generation as an outpainting problem. We consider pixels as rays to outpaint dynamic scenes with the ray context for 3D consistent motions. Experiments demonstrate that our model can generate unbounded dynamic scenes with consistent motions and control the generated contents with scene prompts.
Limitations. Handling reflections, shadow segmentations (shadows of the cartoon man are considered as backgrounds), smaller structure segmentations, full multiview object rendering, complex motions and depth disconnection are exciting future directions.
Acknowledgments. We sincerely thank Prof. Kostas Daniilidis, Dr. Jiahui Lei, Qiao Feng, Chen Wang, and Uday Kiran for their generous and insightful feedback on this work.

### v2

@@ 3.3 Dynamic Scene Update
Point cloud merging. We exploit the depth model [37] to obtain the video depth maps ${\bm{D}}^{(1)}$ of the outpainted video ${\bm{I}}^{(1)}$. Noting that there is a depth inconsistency problem between the estimated depths and the rendered ray depth map ${\bm{\hat{D}}}^{(1)}$ in the mask regions ${\bm{\hat{M}}}^{(1)}$, we follow common practices [58, 57] to align the video depth maps ${\bm{D}}^{(1)}$ with the ray depth maps ${\bm{\hat{D}}}^{(1)}$ by finetuning the depth estimation model. After obtaining the aligned depth maps from the finetuned depth model, we follow §3.1 to initialize foreground point cloud ${\mathcal{P}}^{\mathrm{(1,f)}}$ and background point cloud ${\mathcal{P}}^{\mathrm{(1,b)}}$. Then we merge the new point clouds with the previous one to update the dynamic scenes,
@@
where ${\mathcal{P}}^{(1)}$ is the updated point cloud at the pose $\bm{\Pi}^{(1)}$. In this way, we could generate perpetual dynamic scenes along fly-through camera trajectories by looping §3.2 and §3.3 with new camera poses and scene prompts.
Background completion. In practice, we find that the missing background point clouds occluded by the moving foreground in the scene cause inconsistency when rendering along fly-through camera trajectories (as shown in Figure 4). We hence propose to employ our video outpainting model to generate background videos of the occluded parts. After that, we extract depth maps of the background videos and reconstruct the background point clouds with the extracted depth maps. Details are in supplementary material. We also update the scene with these point clouds.
@@ 4 Experiments
In this section, we first describe the details of our experimental setup, we then discuss the results of the scene generation with fly-through cameras and controllable scene generation. Finally, we conduct ablation studies of our model.
@@ 4.1 Experiment Setups
We trained our outpainting model with the OpenVid dataset [33], which contains high-resolution videos with detailed video prompts. In order to train with daily dynamic scene videos, we sampled 5,000 videos that contain outdoor dynamic scenes, including natural scenes (such as waterfalls, rivers and moving clouds) and urban scenes (such as walking people and moving cars). Additional data curation details can be found in the supplementary material.
@@
Training details. We employed the LoRA [16] to finetune the CogvideoX-5B-I2V model [54]. The LoRA rank was 256256256 and the learning rate was $1\times 10^{-4}$ with a cosine schedule. The video resolution is 16$\times$512$\times$512, and hence $N=16$ and $h=w=512$. For calculating the ray information, we used the depth model [37] to extract the depth map of each video frame. Then, we followed the state-of-the-art video outpainting model MOTIA [46] and exploited the same mask proportion for outpainting.
Camera details. To set up dynamic scenes with fly-through camera trajectories, we followed the common practices in [58, 35, 21], where cameras move along a straight line or rotate. In the former case, we moved the camera backward by 0.0005 units between two adjacent camera poses, where one unit corresponds to the normalized 3D coordinate defined in PyTorch3D, and in the latter case, we rotated the camera by 0.45 radians. We generated camera paths by linearly interpolating translation and approximating the rotation using uniform angular steps.
@@ 4.2 Scene Generation with Fly-through Cameras
Figure 6 shows the dynamic scene generation results with long fly-through trajectories. We tested our model on both the background dynamics (water flowing in the waterfall) and the foreground dynamics (person walking on the path). Compared to the initial point cloud, our model generated perpetual dynamic scenes by exploiting the proposed scene outpainting process. As there is no direct baseline for perpetual dynamic scene generation, we compare our method with the perpetual static scene generation method WonderJourney [58] in Figure 7, state-of-the-art  4D scene generation model CAT4D [49] in Figure 2 and camera-controllable video diffusion model MotionCtrl [48] in Figure 8. These figures demonstrate that our model can better generate perpetual dynamic scenes with 3D consistent motions by exploiting our ray outpainting model.
@@ 4.3 Scene Generation Controlled by Text
To test the controllable dynamic scene generation ability of our model, we generated dynamic scenes from village and volcano images with scene prompts. As shown in Figure 9, for the village image, we control the outpainting content by introducing two different text prompts: waterfall and huge lake. For the volcano image, we prompted the model to generate glacier and magma. Our model successfully controls the perpetual dynamic scene generation with the given scene prompts. Following the common practices in [56, 21, 58, 57], we employed the CLIP [36] scores between the novel view images and the input scene prompts and the dynamic degree (DD) and factual consistency (FC) indices proposed by VideoScore [12] to thoroughly evaluate the generation quality of the video. Table 1 shows that, compared to WonderJourney, our model achieves better performance by controlling dynamic scene generation.
@@ 4.4 Ablation Study
Ray outpainting. To fairly evaluate the proposed ray outpainting model, we conducted experiments on five scenes containing common dynamic objects: cars, people, water and clouds. For each scene, we sampled 16 novel view images from the generated dynamic point cloud at 16 different timestamps along the fly-through camera trajectories. As there are no 3D video outpainting baselines, we chose the 2D outpainting model MOTIA [46] for comparisons. Figure 5 shows that by introducing the ray information for outpainting, our model successfully outpaints 3D consistent motions at the outpainting boundaries. In this way, our ray outpainting model achieves better performance in Table 2.
@@ 5 Conclusion
We proposed DynamicVoyager, a new model for perpetual dynamic scene generation, which reformulates this task as an outpainting problem. By considering pixels as rays and conditioning on ray depth and distance maps, DynamicVoyager can generate scenes with 3D consistent motions. Experimental results demonstrate the superiority of our model.
Limitations. Challenges remain in handling reflections, shadows, fine structures, multi-view rendering, and depth discontinuities, which can be explored in future work.
Acknowledgments
The authors sincerely thank Prof. Kostas Daniilidis, Dr. Jiahui Lei, Qiao Feng, Chen Wang, Ziqing Xu, and Uday Kiran for their generous and insightful feedback on this work. We also acknowledge the support of a Penn Engineering graduate fellowship, Penn startup funds, and the Research Collaboration on the Mathematical and Scientific Foundations of Deep Learning under grants NSF 2031985 and Simons 814201.